OpenStudy (anonymous):

I need to find out how much 6 2/3 cups of green beans in pounds. Its .99 per pound.

OpenStudy (anonymous):

how many oz in cup?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

ok how many oz is 1lb

OpenStudy (anonymous):

16 ounces

OpenStudy (anonymous):

so that means 2cups = 1lb?

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

\[\frac{ x }{ 20/3 }=\frac{ 1 }{ 2 }\] cross multiply

OpenStudy (anonymous):

2x/6.66

OpenStudy (anonymous):

so 3.3 pounds

OpenStudy (anonymous):

2x = 6.6666 x = 6.666/2 = 3.333

OpenStudy (anonymous):

r u cooking or doing math?

OpenStudy (anonymous):

Lol Im trying to figure out how much food and how much my dinners are going to cost for Thanksgiving so in a way both.

OpenStudy (anonymous):

no problem; either u send me 5 bucks or a plate of food

OpenStudy (anonymous):

so the green beans are going to cost 4.32 for 6 2/3 cups

OpenStudy (anonymous):

lol